Suppose that a company sells each unit of its product for $100. The company incurs variable product costs amounting to $30 and variable period costs amounting to $20 for each unit sold. Fixed costs are $300,000. The company desires an after-tax profit of $200,000 and the company’s tax rate is 20%. How many units of its product does the company need to sell to achieve this income target?
Answer #1

After tax target profit = 200,000/80% = 250,000

Units needed to be sold = (Fixed cost+Target profit) /Contribution margin per unit

= (300,000+250,000) / (100-30-20)

= 550,000/50

= 11,000 units